# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: Present address: Nord Quantique, Sherbrooke, QC J1J 2E2, Canada

### v1

@@ Abstract
Dissipative tunneling remains a cornerstone effect in quantum mechanics. In chemistry, it plays a crucial role in governing the rates of chemical reactions, often modeled as the motion along the reaction coordinate from one potential well to another. The relative positions of energy levels in these wells strongly influences the reaction dynamics. Chemical research will benefit from a fully controllable, asymmetric double-well equipped with precise measurement capabilities of the tunneling rates. In this paper, we show that a continuously driven Kerr parametric oscillator with a third order non-linearity can be operated in the quantum regime to create a fully tunable asymmetric double-well. Our experiment leverages a low-noise, all-microwave control system with a high-efficiency readout of the which-well information. We explore the reaction rates across the landscape of tunneling resonances in parameter space. We uncover two new and counter-intuitive effects: (i) a weak asymmetry can significantly decrease the activation rates, even though the well in which the system is initialized is made shallower, and (ii) the width of the tunneling resonances alternates between narrow and broad lines as a function of the well depth and asymmetry. We predict by numerical simulations that both effects will also manifest themselves in ordinary chemical double-well systems in the quantum regime. Our work paves the way for analog molecule simulators based on quantum superconducting circuits.
@@ I Introduction
Engineering Hamiltonians to produce a desired potential landscape is a crucial task in quantum computing [1, 2, 3, 4, 5]. Among these landscapes, double-wells hold particular importance, serving as models for diverse systems like two-level defects [6, 7], nuclear structures [8], and chemical reactions [9, 5]. However, tuning parameters experimentally within these systems, like the barrier height, often proves to be challenging [5, 10]. Additionally, classical computational models can struggle with accuracy, e.g. failing to reach chemical accuracy [11]. Consequently, developing a low-noise system capable of generating tunable double-well potentials is highly desirable for the simulation of quantum chemistry problems.
In this manuscript, we report the results of the activation dynamics of a continuously tunable asymmetric double-well parametric oscillator. During our exploration we found two unexpected effects. First, we find that the asymmetric double-well can experience a significantly longer activation time (well-switching time) from one well to the other than the symmetric one, even when the system is initialized in the shallower well. This is counterintuitive, because one would think that by reducing the barrier height, the activation time should decrease [12, 13, 14]. We find this is not always the case in our system due to a subtle quantum effect described below, which suggests a technique to stabilize bosonic quantum states. The second unexpected effect is that the activation exhibits pronounced quantum resonances whose width alternates between narrow and broad with both the depth and the asymmetry of the wells. This is a manifestation of the width of the Hamiltonian anti-crossing of the energy levels close to the top of the barrier of the double-well energy surface. The location and width of these resonances are well explained within the rotating wave approximation (RWA) and by a semiclassical model, but we note that there is only qualitative agreement between the measured rates and the model.
Based on the experimental and theoretical observation of these effects in the Kerr parametric oscillator (KPO), we investigate whether they are present in other double-well systems. A particularly important class of double-well problems are found in chemistry, for example for modelling electron-transfer reactions [5]. We predict the effects should be generically observable in quantum dissipative double-wells.
@@ II Setup and Model system
Figure 1: Experimental set-up. A) Rendering of the half-aluminum, half-copper sample package containing two sapphire chips magnified in B). Each chip consist of a SNAIL-transmon, a readout resonator, and a Purcell filter. Only one chip is used in this work. The normalized resonance frequency of the SNAIL-transmon is noted $\omega_{a}$. Applying a strong microwave drive at $\omega_{2}\approx 2\omega_{a}$ morphs the SNAIL-transmon Hamiltonian into the parametric oscillator Hamiltonian. C) Schematic of the SNAIL-transmon: a two-SNAIL array serves as the nonlinear element. The capacitor pads are shifted with respect to the axis of the array to couple it to the readout resonator. D) Scanning electron micrograph of the two-SNAIL array. The SNAIL loops are biased with an external magnetic flux $\Phi/\Phi_{0}=0.31$, where $\Phi_{0}$ is the magnetic flux quantum. Figure reproduced from [15].
Our setup consists of two chips with superconducting circuits, shown in Fig. 1 A, that are addressable by microwave drives via charge-coupling. In this experiment, we only make use of one of the two chips. The relevant chip contains an array of two superconducting nonlinear asymmetric inductive elements (SNAILs) [16, 17] shunted by a large capacitor [18], as depicted in Fig. 1 B-D.
@@
where $\omega_{o}$ is the bare resonance frequency of the SNAIL transmon, $g_{3},$ $g_{4}$ are the third- and fourth-order non-linearities of the circuit and $\hat{a}$ is the bosonic annihilation operator. This Hamiltonian is the so-called (asymmetric) parametric oscillator Hamiltonian when $\omega_{2}\approx 2\omega_{o}$ and $\omega_{1}=\omega_{2}/2$ [19, 20]. Here, $\Omega_{1}$ is the amplitude and $\omega_{1}$ the frequency of the drive that will henceforth be referred to as the linear (additive) drive, while $\Omega_{2}$ and $\omega_{2}$ are the amplitude and frequency of what we refer to as the squeezing, or two-photon, parametric drive. The phase $\phi$ is the relative phase between the two drives. By applying displaced frame transformations, transforming into the rotating frame $\omega_{2}/2$ and keeping some terms beyond the RWA [15, 21], we arrive at the effective Hamiltonian describing the asymmetric parametric oscillator
$\frac{\hat{H}_{\text{eff}}}{\hbar}=-K\hat{a}^{\dagger 2}\hat{a}^{2}+\epsilon_{2}(\hat{a}^{2}+\hat{a}^{\dagger 2})+|\epsilon_{1}|(e^{i\phi}\hat{a}+e^{-i\phi}\hat{a}^{\dagger})$ (2)
where $K=-\frac{3g_{4}}{2}+\frac{10g_{3}^{2}}{3\omega_{a}}$ is the leading order Kerr non-linearity, and $\omega_{2}=2\omega_{a}$ with $\omega_{a}\approx\omega_{o}$ the renormalized SNAIL transmon resonance frequency. The drive coefficients are given by $|\epsilon_{1}|=\frac{\Omega_{1}}{2}$ and $\epsilon_{2}=g_{3}\frac{4\Omega_{2}}{3\omega_{a}}$.
Figure 2: Symmetric and asymmetric double-well spectrum. A) Effective double well potential energy $V(x)$, represented here with the quantum energy levels. B) Transition spectrum $|E_{n}-E_{0}|/K$ of the parametric oscillator Hamiltonian as a function of $\epsilon_{2}$ controlling the barrier height [22]. The levels highlighted in green become degenerate at $\epsilon_{2}/K\approx 12$, while the next pair of levels (highlighted in orange) is not yet degenerate at this value of $\epsilon_{2}/K$. C) Same as but A with asymmetry. D) The transition spectrum as a function of $\epsilon_{1}$ controlling the asymmetry. The ground level in the shallow well is highlighted in pink. The pair of levels at the barrier top is highlighted in violet. Note the oscillation of the energy separation between these two levels as a function of $\epsilon_{1}/K$.
@@
where $\hat{a}\mapsto\frac{1}{\sqrt{2}}(x+ip)$ together with $k_{1}=\sqrt{2}|\epsilon_{1}|\cos\phi$, $k_{2}=-\epsilon_{2}$, and $k_{4}=-K/4$. Two instances of $V(x)$ are shown in Fig. 2 A and C, while in Fig. 2 B and D, the associated energy spectra of Eq. 2 are shown as a function of the control parameters $\epsilon_{1}$ and $\epsilon_{2}$ for $\phi=0$. From Eq. 2 or Eq. 3, $\epsilon_{1}$ controls the asymmetry of the wells and $\epsilon_{2}$ controls their depth [19]. Note, however, that the parametric oscillator Hamiltonian cannot be written as a sum of kinetic [$T(p)$] and potential [$V(x)$] energy since cross terms like $x^{2}p^{2}$ are present. These terms can lead to interesting consequences [13, 23, 24, 25] but they do not play a critical role in the new effects described in this paper (see Section IV).
For $\epsilon_{1}=0$, we recover the conventional (symmetric) parametric oscillator Hamiltonian that creates a double-well along the position axis [22, 18, 15, 24, 25, 26]. If $\epsilon_{1}\neq 0$, the phase $\phi$ becomes relevant. For $\phi=90\degree$, the linear drive adds a term proportional to the momentum, $p$, thus not breaking the symmetry of the double-well. For $\phi=0$, this drive adds a term proportional to the position, $x$, thus lifting the degeneracy between the two wells [19, 27] (see Fig. 2 C and D). In the main text, we focus on the case $\phi=0$ and leave the experimental study of the effect of the phase-variation for the Supplementary Material (SM).
@@ III Experiment and analysis
Figure 4: Analysis of the dissipative tunneling resonances according to different models and the experiment. A) Measurement of dissipative tunneling time $T$ as a function of $\epsilon_{2}/K$ (controlling the barrier height) and $\epsilon_{1}/K$ (controlling the asymmetry) in linear color scale. B) Theory prediction from a Lindbladian model including single photon loss and gain with logarithmic color scale. The wavy green-black lines show the maxima of $T$ ‘avoid’ the resonances marked by the red circles. C) Hamiltonian prediction for the resonance conditions (parabolic dashed lines) and EBK’s orbit quantization condition for $n$ and $m$ allowed quantum numbers in the small (orange) and large (blue) figure-8 lobes. The red circles mark triple intersections labeled $(n,m)$.
@@
To measure the activation rate of our system, the states localized at the bottom of the wells need to be prepared and monitored as a function of time. A number of steps are required for this. To ready the setup for our experiments, we bias our SNAIL loops with an external magnetic field sourced by a solenoid lying below the copper part of the enclosure (orange block in Fig. 1 A). This flux allows us to set the Hamiltonian parameters $\omega_{a}$ and $K$. In this work, we choose a flux point at which we directly measure $\omega_{a}/2\pi=6.086$ GHz and $K/2\pi=(528\pm 8)$ kHz (see SM for this and other calibrations). From the flux dependence of both $\omega_{a}$ and $K$ we can fit a model for the SNAIL [17] to extract $g_{3}/3=2\pi\times(-5.6)$ MHz and $g_{4}/4=2\pi\times(-74)$ kHz. The values of $\Omega_{1}$ and $\Omega_{2}$ are directly proportional to the microwave amplitude we apply to our sample and, therefore, we have precise control over $\epsilon_{1}$ and $\epsilon_{2}$. In our system, dielectric loss sets the single-photon lifetime to 20 µs.
@@
By turning on the squeezing drive, and waiting five times the single-photon lifetime, we prepare the oscillator in its steady state. We observe the bifurcation of our SNAIL oscillator by homodyning the emitted radiation activated by a tone parametrically coupling the parametric oscillator with the on-chip readout resonator itself coupled to our quantum-limited amplifier detection line [18]. The homodyne signal clearly shows the typical pair of stable oscillations out-of-phase by $180^{\circ}$. Importantly, the photons emitted by the oscillator during readout are continuously replenished by the squeezing drive: the driven oscillator in presence of dissipation remains in one of its two quasi-steady states. Acquiring a single data point for the which-well information takes 4 µs, which is typically much shorter than the activation time across the double-well parametric oscillator barrier. We postselect the instances when the system is initially found in the shallower well. By performing a second measurement after a variable waiting time (see SM), one can detect well-switching. By repeating these measurements, one can determine the probability per unit time of witnessing an activation event and therefore obtain the activation rate for a given set of Hamiltonian parameters $\epsilon_{1}/K$ and $\epsilon_{2}/K$.
@@
In Fig. 3 A we show measurements of the population dynamics of the wells for different values of the asymmetry, controlled by $\epsilon_{1}/K$. For these measurements the squeezing amplitude is set to $\epsilon_{2}/K=7.7$. The probability, as a function of time, of being in the initial well is well approximated by an exponential decay (see SM). This timescale $T$ of this exponential is a direct measurement of the activation rate ($1/T$). In Fig. 3 B we see a theory prediction computed from a Lindbladian model including single-photon loss at rate $\kappa/K=0.025$ and gain associated with a finite temperature corresponding to a mean number of thermal photons $n_{\mathrm{th}}=0.05$ (or equivalently 95 mK at 6 GHz). The agreement between theory and experiment for the variations of $T$ with well parameters is remarkable, except when we consider the measured scale of $T$, which differs from theory by more than an order of magnitude. This remains only qualitatively understood [30].
@@
We observe in both the theory and in the experiment resonances for certain values of $\epsilon_{1}/K$, where the activation rate is markedly increased. These are resonances between levels localized in different wells. The resonances for levels deep within the wells behave effectively as level crossings since the coupling is exponentially small due to the suppression of tunneling under the barrier. Therefore, these activation events are mediated by thermal and quantum heating [14, 13] from the ground state into the tunneling levels at the barrier top.
@@
where $\tilde{n}=n,m$ are the odd quantum numbers of the shallow and deeper wells respectively. The triple intersection points of the parabolas from Eq. 4 with the equi-action curves meeting the quantization condition mark the point in parameter space where a new level enters the wells in the tunneling resonances condition. We mark them with red circles labeled by $(n,m)$, the quantum number of each well. At these points, where each well contains exactly $(n+1)/2$ and $(m+1)/2$ semiclassical orbits, we expect the resonance to broaden due to a new orbit contributing to the activation rate. This is seen as sharpening structures in Fig. 4 A and B (see also Fig. 3 for the broadened resonances).
@@
In Fig. 4 B, we show the Lindbladian model of the experiment. The agreement is excellent with regard to the location and behavior of the resonance in parameter space, but falls short in quantitatively predicting the activation rates measured. However, many features of the data are correctly captured. For example, the Lindbladian model shows that an asymmetric system can have a longer activation lifetime than the symmetric system, even if one of the wells is markedly shallower. See this directly from Fig. 3 at $\epsilon_{1}/K\approx 1$. The green-black line in Fig. 4 A shows the experimentally determined maxima of activation time $T$ as a function of the control parameters. The theoretical maxima are shown in Fig. 4 B by the line of the same colors. The location of the maxima is nontrivial, and it is relevant because it can readily be used to extend the lifetime of a Kerr-cat qubit [18]. The effect was unknown to us before the analysis of our experimental data.
Figure 5: Experimental and theoretical activation time $T$ for symmetric KPO and at optimal asymmetry. Depicted as function of barrier height $\epsilon_{2}/K$. A-B) The symmetric case is modulated by resonances in the quantized energies (see [15] and SM). Exploiting asymmetry, it is possible to avoid these resonances and substantially increase the activation time. The asymmetric optimum corresponds to the green-black lines in Fig. 4 A, B.
In Fig. 5, we show the increase of $T$ with the asymmetry. To provide physical insight into this effect, we note that for the symmetric case ($\epsilon_{1}=0$), $T$ is modulated in a step-like fashion by the orbits falling under the barrier [15]. The linear drive can be exploited to break the (parity) symmetry of these orbits as shown in Fig. 2 and therefore avoid altogether the resonant saturation seen for $\epsilon_{1}=0$ in Fig. 5 A, B. This observation explains the trajectory of the green-black curve in Fig. 4 A, B which snakes around, avoiding the new resonance conditions marked by the red circles. In other words, quantum tunneling produces a hybridization of the classically decoupled orbits right under the barrier, so the asymmetry parameter can be used to minimize that hybridization, reducing the tunneling rate via the excited states, and avoid the plateaus in Fig. 5 almost completely.

### v2

@@ Abstract
Dissipative tunneling remains a cornerstone effect in quantum mechanics. In chemistry, it plays a crucial role in governing the rates of chemical reactions, often modeled as the motion along the reaction coordinate from one potential well to another. The relative positions of energy levels in these wells strongly influence the reaction dynamics. Chemical research will benefit from a fully adjustable, asymmetric double-well equipped with precise measurement capabilities of the tunneling rates. In this paper, we show a quantum simulator system that consists of a continuously driven Kerr parametric oscillator with a third order non-linearity that can be operated in the quantum regime to create a fully tunable asymmetric double-well. Our experiment leverages a low-noise, all-microwave control system with a high-efficiency readout, based on a tunnel Josephson junction circuit, of the which-well information. We explore the reaction rates across the landscape of tunneling resonances in parameter space. We uncover two new and counter-intuitive effects: (i) a weak asymmetry can significantly decrease the activation rates, even though the well in which the system is initialized is made shallower, and (ii) the width of the tunneling resonances alternates between narrow and broad lines as a function of the well depth and asymmetry. We predict by numerical simulations that both effects will also manifest themselves in ordinary chemical double-well systems in the quantum regime. Our work is a first step for the development of analog molecule simulators of proton transfer reactions based on quantum superconducting circuits.
@@ I Introduction
Engineering Hamiltonians to produce a desired potential landscape is a crucial task in quantum computing [1, 2, 3, 4, 5, 6]. Among these landscapes, double-wells hold particular importance, serving as models for diverse systems like two-level defects [7, 8], nuclear structures [9], and chemical reactions [10, 6]. However, tuning parameters experimentally within these systems, like the barrier height, often proves to be challenging [6, 11]. Additionally, classical computational models can struggle with accuracy, e.g. failing to reach chemical accuracy [12]. Consequently, developing a low-noise system capable of generating tunable double-well potentials is highly desirable for the simulation of quantum chemistry problems.
In this manuscript, we report the results of the activation dynamics of a continuously tunable asymmetric double-well parametric oscillator. During our exploration, we found two unexpected effects. First, we find that the asymmetric double-well can experience a significantly longer activation time (well-switching time) from one well to the other than the symmetric one, even when the system is initialized in the shallower well. This is counterintuitive, because one would think that by reducing the barrier height, the activation time should decrease [13, 14, 15]. We find this is not always the case in our system due to a subtle quantum effect described below, which suggests a technique to stabilize bosonic quantum states. The second unexpected effect is that the activation exhibits pronounced quantum resonances whose width alternates between narrow and broad with both the depth and the asymmetry of the wells. This is a manifestation of the width of the Hamiltonian anti-crossing of the energy levels close to the top of the barrier of the double-well energy surface. The location and width of these resonances are well explained by a Hamiltonian model within the rotating wave approximation (RWA) and by a semiclassical model, but we note that there is only qualitative agreement between the measured rates and a simple model of dissipation.
Based on the experimental and theoretical observation of these effects in the Kerr parametric oscillator (KPO), we investigate whether they are present in other double-well systems. A particularly important class of double-well problems are found in chemistry, for example for modeling electron-transfer reactions [6]. We predict the effects should be generically observable in quantum dissipative double-wells.
@@ II Setup and Model system
Figure 1: Experimental set-up. A) Rendering of the half-aluminum, half-copper sample package containing two sapphire chips magnified in B). Each chip consists of a SNAIL-transmon, a readout resonator, and a Purcell filter. Only one chip is used in this work. The normalized resonance frequency of the SNAIL-transmon is noted $\omega_{a}$. Applying a strong microwave drive at $\omega_{2}\approx 2\omega_{a}$ morphs the SNAIL-transmon Hamiltonian into the parametric oscillator Hamiltonian. C) Schematic of the SNAIL-transmon: a two-SNAIL array serves as the nonlinear element. The capacitor pads are shifted with respect to the axis of the array to couple it to the readout resonator. D) Scanning electron micrograph of the two-SNAIL array. The SNAIL loops are biased with an external magnetic flux $\Phi/\Phi_{0}=0.31$, where $\Phi_{0}$ is the magnetic flux quantum. Figure adapted from [16].
Our setup was first described in [17], with the current implementation first introduced in [16] and summarized here for the sake of completeness. The setup consists of two chips with superconducting circuits, shown in Fig. 1 A, that are addressable by microwave drives via charge-coupling. In this experiment, we only make use of one of the two chips. The relevant chip contains an array of two superconducting nonlinear asymmetric inductive elements (SNAILs) [18, 19] shunted by a large capacitor [17], as depicted in Fig. 1 B-D.
@@
where $\omega_{o}$ is the bare resonance frequency of the SNAIL transmon, $g_{3},$ $g_{4}$ are the third- and fourth-order non-linearities of the circuit and $\hat{a}$ is the bosonic annihilation operator. This Hamiltonian is the so-called (asymmetric) parametric oscillator Hamiltonian when $\omega_{2}\approx 2\omega_{o}$ and $\omega_{1}=\omega_{2}/2$ [20, 21]. Here, $\Omega_{1}$ is the amplitude and $\omega_{1}$ the frequency of the one-photon drive that will henceforth be referred to as the linear, or additive, drive, while $\Omega_{2}$ and $\omega_{2}$ are the amplitude and frequency of what we refer to as the squeezing, or two-photon, parametric drive. The phase $\phi$ is the relative phase between the two drives. By applying displaced frame transformations, transforming into the rotating frame at $\omega_{2}/2$ and keeping some terms beyond the RWA [16, 22], we arrive at the effective Hamiltonian describing the asymmetric parametric oscillator
$\frac{\hat{H}_{\text{eff}}}{\hbar}=-K\hat{a}^{\dagger 2}\hat{a}^{2}+\epsilon_{2}(\hat{a}^{2}+\hat{a}^{\dagger 2})+\epsilon_{1}(e^{i\phi}\hat{a}+e^{-i\phi}\hat{a}^{\dagger})$ (2)
where $K=-\frac{3g_{4}}{2}+\frac{10g_{3}^{2}}{3\omega_{a}}$ is the leading order Kerr non-linearity, and $\omega_{2}=2\omega_{a}$ with $\omega_{a}\approx\omega_{o}$ the renormalized SNAIL transmon resonance frequency. The drive coefficients are given by $\epsilon_{1}=\frac{\Omega_{1}}{2}$ and $\epsilon_{2}=g_{3}\frac{4\Omega_{2}}{3\omega_{a}}$.
Figure 2: Symmetric and asymmetric double-well spectrum. A) Effective double well potential energy $V(x)$, represented here with the quantum energy levels. B) Transition spectrum $|E_{n}-E_{0}|/K$ of the parametric oscillator Hamiltonian as a function of $\epsilon_{2}$ controlling the barrier height [23]. The levels highlighted in green become degenerate at $\epsilon_{2}/K\approx 12$, while the next pair of levels (highlighted in orange) is not yet degenerate at this value of $\epsilon_{2}/K$. C) Same as A but with asymmetry. D) The transition spectrum as a function of $\epsilon_{1}$ controlling the asymmetry. The ground level in the shallow well is highlighted in pink. The pair of levels at the barrier top is highlighted in violet. Note the oscillation of the energy separation between these two levels as a function of $\epsilon_{1}/K$.
@@
where $\hat{a}\mapsto\frac{1}{\sqrt{2}}(x+ip)$ together with $k_{1}=\sqrt{2}\epsilon_{1}\cos\phi$, $k_{2}=-\epsilon_{2}$, and $k_{4}=-K/4$. Two instances of $V(x)$ are shown in Fig. 2 A and C, while in Fig. 2 B and D, the associated energy spectra of Eq. 2 are shown as a function of the control parameters $\epsilon_{1}$ and $\epsilon_{2}$ for $\phi=0$. From Eq. 2 or Eq. 3, $\epsilon_{1}$ controls the asymmetry of the wells and $\epsilon_{2}$ controls their depth [20]. Note, however, that the parametric oscillator Hamiltonian cannot be written as a sum of kinetic [$T(p)$] and potential [$V(x)$] energy since cross-terms like $x^{2}p^{2}$ are present. These terms can lead to interesting consequences [14, 24, 25, 26], but they do not play a critical role in the new effects described in this paper (see Section IV).
For $\epsilon_{1}=0$, we recover the conventional (symmetric) parametric oscillator Hamiltonian that creates a double-well along the position axis [23, 17, 16, 25, 26, 27]. If $\epsilon_{1}\neq 0$, the phase $\phi$ becomes relevant. For $\phi=90\degree$, the linear drive adds a term proportional to the ‘momentum’, $p$, thus not breaking the symmetry between the wells. For $\phi=0$, this drive adds a term proportional to the ‘position’, $x$, thus lifting the degeneracy between the two wells [20, 28] (see Fig. 2 C and D). In the main text, we focus on the case $\phi=0$ and leave the experimental study of the effect of the phase-variation for the Supplementary Material (SM).
@@ III Experiment and analysis
Figure 4: Analysis of the dissipative tunneling resonances according to different models and the experiment. A) Measurement of dissipative tunneling time $T$ as a function of $\epsilon_{2}/K$ (controlling the barrier height) and $\epsilon_{1}/K$ (controlling the asymmetry) in linear color scale. B) Theory prediction from a Lindbladian model including single photon loss and gain with logarithmic color scale. The wavy green-black lines show that the maxima of $T$ ‘avoid’ the resonances marked by the red circles. C) Hamiltonian prediction for the resonance conditions (parabolic dashed lines) and EBK’s orbit quantization condition for $n$ and $m$ allowed quantum numbers in the small (orange) and large (blue) figure-8 lobes. The red circles mark triple intersections labeled $(n,m)$.
@@
To measure the activation rate of our system, the states localized at the bottom of the wells need to be prepared and monitored as a function of time. A number of steps are required for this. To ready the setup for our experiments, we bias our SNAIL loops with an external magnetic field sourced by a solenoid lying below the copper part of the enclosure (orange block in Fig. 1 A). This flux allows us to set the Hamiltonian parameters $\omega_{a}$ and $K$. In this work, we choose a flux point at which we directly measure $\omega_{a}/2\pi=6.086$ GHz and $K/2\pi=(528\pm 10)$ kHz (see SM for this and other calibrations). From the flux dependence of both $\omega_{a}$ and $K$ we can fit a model for the SNAIL [19] to extract $g_{3}/3=2\pi\times(-5.6)$ MHz and $g_{4}/4=2\pi\times(-74)$ kHz. The values of $\Omega_{1}$ and $\Omega_{2}$ are directly proportional to the microwave amplitude we apply to our sample and, therefore, we have precise control over $\epsilon_{1}$ and $\epsilon_{2}$. In our system, dielectric loss sets the single-photon lifetime to 20 µs.
@@
By turning on the squeezing drive, and waiting five times the single-photon lifetime, we prepare the oscillator in its steady state. We observe the bifurcation of our SNAIL oscillator by homodyning the emitted radiation activated by a tone parametrically coupling the parametric oscillator with the on-chip readout resonator, itself coupled to our quantum-limited amplifier detection line [17]. The homodyne signal clearly shows the typical pair of stable oscillations out-of-phase by $180^{\circ}$. Importantly, the photons emitted by the oscillator during readout are continuously replenished by the squeezing drive: the driven oscillator in presence of dissipation remains in one of its two quasi-steady states. Acquiring a single data point for the which-well information takes 4 µs, which is typically much shorter than the activation time across the double-well parametric oscillator barrier, which is of order 100 µs. We postselect the instances when the system is initially found in the shallower well. By performing a second measurement after a variable waiting time (see SM), one can detect well-switching. By repeating these measurements, one can determine the probability per unit time of witnessing an activation event and therefore obtain the activation rate for a given set of Hamiltonian parameters $\epsilon_{1}/K$ and $\epsilon_{2}/K$.
@@
In Fig. 3 A we show measurements of the population dynamics of the wells for different values of the asymmetry, controlled by $\epsilon_{1}/K$. For these measurements the squeezing amplitude is set to $\epsilon_{2}/K=7.7$. The probability, as a function of time, of being in the initial well is well approximated by an exponential decay (see SM). This timescale $T$ of this exponential is a direct measurement of the activation rate ($1/T$). In Fig. 3 B we see a theory prediction computed from a Lindbladian model including single-photon loss at rate $\kappa/K=0.025$ and gain associated with a finite temperature corresponding to a mean number of thermal photons $n_{\mathrm{th}}=0.05$ (or equivalently 95 mK at 6 GHz). The agreement between theory and experiment for the variations of $T$ with well parameters is remarkable, except when we consider the measured scale of $T$, which differs from theory by more than an order of magnitude. While existing models [31, 27, 16] fail to quantitatively predict this activation time, our findings demonstrate that the underlying physics can be understood without quantitative predictions of $T$.
@@
We observe, in both theory and experiment, resonances for certain values of $\epsilon_{1}/K$, where the activation rate is markedly increased. These are resonances between levels localized in different wells. The resonances for levels deep within the wells behave effectively as level crossings since the coupling is exponentially small due to the suppression of tunneling under the barrier. Therefore, these activation events are mediated by thermal and quantum heating [15, 14] from the ground state into the tunneling levels at the barrier top.
@@
where $\tilde{n}=n,m$ are the quantum numbers of the shallow and deep wells respectively. The triple intersection points of the parabolas from Eq. 4 with the equi-action curves meeting the EBK quantization condition mark the point in parameter space where a new level enters the wells in the tunneling resonances condition. We mark them with red circles labeled by $(n,m)$, the quantum number of each well. At these points, where each well contains exactly $n$ and $m$ semiclassical orbits, we expect the resonance to broaden due to a new orbit contributing to the activation rate. This is seen as sharpening structures in Fig. 4 A and B (see also Fig. 3 for the broadened resonances).
@@
In Fig. 4 B, we show the predictions of the Lindbladian model of the experiment. The agreement is excellent with regards to the location and behavior of the resonance in parameter space, but falls short in quantitatively predicting the activation rates measured. However, many features of the data are correctly captured. For example, the Lindbladian model shows that an asymmetric system can have a longer activation lifetime than the symmetric system, even if one of the wells is markedly shallower (see Fig. 3 at $\epsilon_{1}/K\approx 1$). The green-black dashed line in Fig. 4 A shows the experimentally determined maxima of activation time $T$ as a function of the control parameters. The theoretical maxima are shown in Fig. 4 B by the line with the same colors. The location of the maxima is nontrivial, and it is relevant because it can readily be used to extend the lifetime of a Kerr-cat qubit [17]. The effect was unknown to us before the analysis of our experimental data.
Figure 5: Experimental and theoretical activation time $T$ for symmetric KPO and at optimal asymmetry. Depicted as a function of barrier height $\epsilon_{2}/K$. A-B) The symmetric case is modulated by resonances in the quantized energies (see [16] and SM). Exploiting asymmetry, it is possible to avoid these resonances and substantially increase the activation time. The asymmetric optimum corresponds to the green-black lines in Fig. 4 A, B.
In Fig. 5, we show the increase of $T$ with the asymmetric parameter. To provide physical insight into this effect, we note that for the symmetric case ($\epsilon_{1}=0$, plotted in orange), $T$ is modulated in a step-like fashion as the quantized orbits fall under the barrier [16]. The linear drive can be exploited to break the parity symmetry of these orbits as shown in Fig. 2 and therefore avoid altogether the resonant saturation seen for $\epsilon_{1}=0$ in Fig. 5 A, B. This observation explains the trajectory of the green-black curve in Fig. 4 A, B, which avoids the new resonance conditions marked by the red circles. In other words, quantum tunneling produces a hybridization of the classically decoupled orbits right under the barrier. The asymmetry parameter can be adjusted to minimize that hybridization, reducing the tunneling rate via the excited states, and avoid the plateaus in Fig. 5 almost completely.